Calculate the pH of a solution prepared by dissolving 2.20 got sodium acetate, CH3COONa, in 78.0 mL of 0.25 M acetic acid, CH3COOH(aq). Assume the volume change upon dissolving the sodium acetate is negligible. Ka of CH3COOH is 1.75 x 10^-5


1
Expert's answer
2021-11-13T01:33:22-0500

Moles of CH3COONa "= \frac{2.20}{82.0} = 0.02683 \;mol"

Volume V = 78.0 mL = 0.078 L

[CH3COONa] "= \frac{0.02683}{0.078} = 0.3439 \;M"

[CH3COOH] = 0.25 M

"pH = -log(K_a) + log(\frac{[CH_3COONa]}{[CH_3COOH]}) \\

pH = -log(1.75 \times 10^{-5}) + log(\frac{0.3439}{0.25}) \\

pH = 4.89"